On a right triangle with base measured as x, the side labled as x, and the hypotenuse unlabeled,and the perimeter of the triangle is 4+2square root2. What is the value of x? the answer choices are A) 2 B) 4 C) square root of 2 D) 2Square root2 and E) 2+2square root2. Can you please respond quickley and show work thanks.
Answer by rothauserc(4718)

let h be the hypotenuse
h^2 = x^2 + x^2
h^2 = 2x^2
h = x * square root(2)
*****************************************************
x + x + (x*square root(2) = 4+2square root(2)
x * (2 + square root(2)) = 4 + 2square root(2)
x = (4 + 2square root(2)) / (2 + square root(2))
x = 2(2 + square root(2)) / (2 + square root(2))
x = 2
******************************************************
answer is A) 2